A person wears glasses of power – 2.5 D. Is the person farsighted or nearsighted? What is the far point of the person without the glasses?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. The person wears glasses of power –2.5D So, the person must be near sighted. u = ∞, v = far point, f = 1/-25 = - 0.4m = - 40 cm now, 1/v – 1/u = 1/f ⇒ 1/v = 1/u + 1/f = 0 + 1/-40 ⇒ v = - 40 cm So, the far point of the person is 40 cm